Find the equation of the line containing the point and perpendicular to the given line. (0,–7); 4x + 3 = y please show work.
Answer by Alan3354(69443)

Find the equation of the line containing the point and perpendicular to the given line. (0,–7); 4x + 3 = y
---------
y = 4x + 3 is in slope-intercept form, so slope = 4
Lines perpendicular have a slope that's the negative inverse, m = -1/4
Use y = mx + b and the point to find b
-7 = m*0 + b
b = -7
---------
--> y = (-1/4)x - 7
x + 4y = -28